# On total transitivity of graphs

### v1

@@ 2.2 Basic properties of total transitivity
For any graph $G$, $Tr_{t}(G)\leq min\{\Delta(G),\left\lfloor\frac{n}{2}\right\rfloor\}$, where $\Delta(G)$ is the maximum degree of $G$.
@@
Let $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$ be a $Tr_{t}$-transitive partition of $G$. Let $x\in V_{k}$. As $V_{i+1}$ is a total dominating set in the graph $G_{i}=G-\bigcup_{j=1}^{i}V_{i}$, $deg(x)\geq k$ and $V_{i}\geq 2$ for all $i$. Therefore, $Tr_{t}(G)=k\leq deg(x)\leq\Delta(G)$ and $Tr_{t}(G)\leq\left\lfloor\frac{n}{2}\right\rfloor$. Hence, $Tr_{t}(G)\leq min\{\Delta(G),\left\lfloor\frac{n}{2}\right\rfloor\}$. ∎
@@
The partition $\{V_{1},V_{2},\ldots,V_{\left\lfloor\frac{n}{2}\right\rfloor}\}$ such that each set contains at least two vertices is a total transitive partition of $K_{n}$ of size $\left\lfloor\frac{n}{2}\right\rfloor$. Moreover, from the Proposition 1, we have $Tr_{t}(G)\leq min\{\Delta(K_{n}),\left\lfloor\frac{n}{2}\right\rfloor\}=min\{n,\left\lfloor\frac{n}{2}\right\rfloor\}=\left\lfloor\frac{n}{2}\right\rfloor$. Therefore, $Tr_{t}(K_{n})=\left\lfloor\frac{n}{2}\right\rfloor$. ∎
@@
$Tr_{t}(P_{n})=\begin{cases}1&n=2\\ 2&n\geq 3\end{cases}$
@@
Let $C_{n}$ be a cycle of $n$ vertices, then the total transitivity of $C_{n}$ is $2$.
@@
For a complete bipartite graph $K_{m,n}$ with $m+n$ vertices $Tr_{t}(K_{m,n})=min\{m,n\}$.
@@
Let the vertex set of $K_{m,n}$ be $V=X\cup Y$ such that $|X|=m,|Y|=n$ and $m\leq n$. Consider a vertex partition $\pi=\{V_{1},V_{2},\ldots,V_{m}\}$, where $V_{i}=\{x_{i},y_{i}\}$, for all $2\leq i\leq m$ and $V_{1}$ contains all the other vertices of $K_{m,n}$. Since $K_{m,n}$ is a complete bipartite graph, $\pi$ is a total transitive partition of size $m$ and $Tr_{t}(K_{m,n})\geq m$. Next, we show that $Tr_{t}(K_{m,n})$ cannot be more than $m$. If possible, let $Tr_{t}(K_{m,n})=k\geq m+1$ and $\pi^{\prime}=\{V_{1},V_{2},\ldots,V_{k}\}$ be a total transitive partition of $K_{m,n}$ of size $k$. Let $v\in V_{k}$, then the degree of $v$ must be at least $k\geq m+1$. So, the vertex $v$ cannot be from $Y$. Moreover, as $V_{k}$ dominates $V_{k}$, there exists $u\in V_{k}$ such that $uv\in E$ and $u\in Y$. Since $u\in V_{k}$, the degree of $u$ must be at least $k\geq m+1$. We have a contradiction as $deg(u)=m<k$. Therefore, $Tr_{t}(K_{m,n})\leq m$. Hence, $Tr_{t}(K_{m,n})=min\{m,n\}$. ∎
@@
If $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$ is a total transitive partition of a graph $G$, then for any two sets $V_{i}$ and $V_{j}$ in $\pi$, where $i<j$, the partition $\pi=\{V_{1},V_{2},\ldots,V_{i-1},V_{i}\cup V_{j},V_{i+1},\ldots,V_{j-1},V_{j+1},\ldots,V_{k}\}$ is a total transitive partition.
@@
Let $G$ be a connected graph and $TTr(G)=k$, $k\geq 2$. Then there exists a total transitive partition of $G$ of size $k$, say $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$ of $G$, such that $|V_{k}|=2$.
Now we show that if the total transitivity of a graph $G$ is constant, that is, independent of the number of vertices, and $\pi=\{V_{1},V_{2},\ldots,V_{k-1},V_{k}\}$ is a total transitive partition of $G$, then the number of vertices of each $V_{i}$, for $2\leq i\leq k$, does not depend on $n$, rather it is bounded by some exponential in $i$.
Lemma 9.
@@
Let $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$ be a total transitive partition of $G$ of size $k$, where $k\geq 3$. Let $x\in V_{k}$, as $V_{k}$ dominates $V_{k}$, there exists $y\in V_{k}$ such that $xy\in E(G)$. So, $|V_{k}|\geq 2$. Now if $|V_{k}-\{x,y\}|\geq 1$, then we move all the vertex except two, say vertices $x,y$, from the set $V_{k}$ to the first set $V_{1}$. After this transformation, we have, $\pi^{\prime}=\{V_{1}^{\prime}=V_{1}\cup(V_{k}-\{x,y\}),V_{2},\ldots,V_{k-1},V_{k}^{\prime}=\{x,y\}\}$ which is a total transitive partition of $G$ of size $k$ and $|V_{k}|=2$. Therefore, $\pi^{\prime}$ is a total transitive of $G$ of size $k$, such that $|V_{k}|=2$. Consider the base case when $i=1$. Note that to dominate $V_{k}$ and $V_{k-1}$, we need at most $2\cdot|V_{k}|$ vertices in, $V_{k-1}$ and the remaining vertices can be moved to, $V_{1}$ as in the previous argument. Therefore, $|V_{k-1}|\leq 2\cdot|V_{k}|=2\cdot 2=4=2^{i+1}$. Assume the induction hypothesis and consider that we have a total transitive partition $\pi=\{V_{1},V_{2},\ldots\,V_{k}\}$, such that $|V_{k}|=2$ and $|V_{k-j}|\leq 2^{j+1}$, $11\leq j\leq i-$ 1. Now consider the set $V_{k-i}$. Similarly, to dominate the $i+1$ sets that are following it and including it, we need at most $2\cdot|V_{k-i+1}|$ vertices in $V_{k-i}$. If $V_{k-i}$ contains more than those many vertices, the remaining vertices can be moved to $V_{1}$, maintaining total transitivity of $G$. Hence, $V_{k-i}\leq 2\cdot|V_{k-i+1}|=2\cdot 2^{i}=2^{i+1}$. ∎
When the number of vertices in each set is bounded by a constant, then the number of possible subsets that can be $V_{1}$ is polynomial in $n$, where $n$ is the number of vertices of the graph. The same is true for every set $V_{2},V_{3},\ldots,V_{k}$. Therefore, we have a polynomial (in $n$) number of possible solutions for Maximum Total Transitivity Decision Problem when the total transitivity of the graph is constant. Also, note that we can verify whether a vertex partition is a total transitive partition or not in polynomial time. Therefore, when the total transitivity of the graph is constant, we can solve Maximum Total Transitivity Decision Problem in polynomial time. Hence, we have the following corollary.
Corollary 10.
The Maximum Total Transitivity Decision Problem is a fixed parameter tractable when parameterized by the size of the total transitive partition.
In the Remark 1, we have already shown that the difference between transitivity and total transitivity can be arbitrarily large. Now we show another non-trivial example where the difference between these two parameters is arbitrarily large. Let $G=(S\cup K,E)$ be a split graph, where $S$ and $K$ are an independent set and clique of $G$, respectively. Also, assume that $K$ is the maximum clique of $G$, that is, $\omega(G)=|K|$. Furthermore, $|K|=|S|$, and each vertex of $S$ is adjacent to exactly one vertex of $K$ (see Figure 1).
@@
First, we show that for the above split graph, $Tr_{t}(G)=1$. Let $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$ be a total transitive partition of the above split graph $G$. Since $deg(s)=1$ for all $s\in S$, $s$ cannot be in $V_{p}$ for $p\geq 2$. Moreover, $\pi$ is a total transitive, which implies that to dominate $s\in S$, there must be a vertex from $K$ in $V_{1}$. From the above example graph, it is clear that to dominate vertices from $S$, all the vertices of $K$ must be in $V_{1}$. Hence, $Tr_{t}(G)=1$. Again, we know from [18], $Tr(G)=\omega(G)+1$, as $S$ dominates $K$. Therefore, for large $\omega(G)$, the difference $Tr(G)-Tr_{t}(G)=\omega(G)+1-1=\omega(G)$ is arbitrarily large.
Next, we prove the existence of graphs with transitivity is one more than total transitivity. We define a special tree, namely total complete minimum broadcast tree of order $k$ ($tcmbt_{k}$). Apart from this property, the $tcmbt_{k}$ possesses some important properties. The recursive definition of $tcmbt_{k}$ is as follows:
@@
A total complete minimum broadcast tree of order $k$ $(tcmbt_{k})$ is defined recursively as follows:
@@
    The tcmbt of order $1$ $(tcmbt_{k})$ is a rooted tree, whose root has degree $1$ and adjacent to a vertex of degree $1$.
@@
    The tcmbt of order $k$ $(tcmbt_{k})$ is a rooted tree, whose root has degree $k$ and is adjacent to vertices $\{v_{1},v_{2},\ldots,v_{k}\}$, such that $v_{i}$ are roots of the tcmbt of order $i$ $(tcmbt_{i})$, for all $11\leq i\leq k-$ 1. Moreover, the vertex $v_{k}$ is adjacent to vertices $\{u_{1},u_{2},\ldots,u_{k-1}\}$, such that $u_{i}$ are roots of the tcmbt of order $i$ $(tcmbt_{i})$, for all $11\leq i\leq k-$ 1.
Figure 2: $tcmbt$ of order upto $3$ along with its unique total transitive partition
Figure 2, illustrates the $tcmbt$ of order upto $3$. Note that from the definition, it follows that the root of $tcmbt_{k}$ and exactly one of its children are the only vertices that can be in $V_{k}$, in a total transitive partition of order $k$. Because of the acyclic property, none of the vertices in $V_{i}$ can dominate more than one vertex from $V_{j}$ with $j>i$. Therefore, the $tcmbt_{k}$ is the minimum tree (in terms of vertices) with total transitivity as $k$, and it has a unique total transitive partition of size $k$ (refer Figure 2). This property of $tcmbt_{k}$ helps us to use these graphs as a gadget in the NP-completeness proofs in Section 4. We show that if a graph $G$ is a $tcmbt_{k}$, then $Tr_{t}(G)=k$ and $Tr(G)=k+1$. For that, we have the following lemma.
@@
First, we show that $Tr_{t}(T)=k$. For this, we use induction on the order of $tcmbt$. For the base case, that is, when $k=1$, by the definition of $tcmbt$, $T$ is a $P_{2}$. Therefore, $Tr_{t}(P_{2})=1$. Let $T$ be a $tcmbt_{k}$. By the definition, we have $T$ is a rooted tree, whose root has degree $k$ and is adjacent to vertices $\{v_{1},v_{2},\ldots,v_{k}\}$, such that $v_{i}$ are roots of the tcmbt of order $i$ $(tcmbt_{i})$, for all $11\leq i\leq k-$ 1. Moreover, the vertex $v_{k}$ is adjacent to vertices $\{u_{1},u_{2},\ldots,u_{k-1}\}$, such that $u_{i}$ are roots of the tcmbt of order $i$ $(tcmbt_{i})$, for all $11\leq i\leq k-$ 1. Let $c$ be the root of $T$ and $c^{\prime}$ be its child such that $c^{\prime}$ is adjacent to vertices $\{u_{1},u_{2},\ldots,u_{k-1}\}$, such that $u_{i}$ are roots of the tcmbt of order $i$ $(tcmbt_{i})$, for all $11\leq i\leq k-$ 1. By the induction hypothesis, $Tr_{t}(T_{v_{j}})=Tr_{t}(T_{u_{j}})=j$, for all $11\leq i\leq k-$ 1, where $T_{v}$ is a subtree of $T-\{c\}$, rooted at $v$ (see the $tcmbt_{3}$ in Figure 2). Now $T-\{c,c^{\prime}\}$ is a disjoint union of trees $T_{v_{j}},T_{u_{j}}$, for all $11\leq j\leq k-$ 1, and $k-1=\max\{Tr_{t}(T_{v_{j}})|1\leq j\leq k-1\}=\max\{Tr_{t}(T_{u_{j}})|1\leq j\leq k-1\}$. As total transitivity depends only on the maximum total transitivity of a component of a disconnected graph (PropositionLABEL:disjoint_union_total_transitivity). Therefore, $Tr_{t}(T-\{c,c^{\prime}\})$ is $k-1$. Moreover, we can construct a total transitive partition, say $\pi=\{V_{1},V_{2},\ldots,V_{k-1}\}$, of $T-\{c,c^{\prime}\}$ recursively such that $\{v_{i},u_{i}\}\in V_{i}$ for all $11\leq i\leq k-$ 1. Now clearly, $\pi^{\prime}=\pi\cup\{c,c^{\prime}\}$ forms a total transitive partition of $T$. Therefore, $Tr_{t}(T)\geq k$. Moreover, by the Proposition 1, we know that $Tr_{(}T)\leq\Delta(G)=k$. Hence, $Tr_{t}(T)=k$. The algorithm described in [14] shows that $Tr(T)=k+1$. Therefore, $Tr_{t}(T)=k$ and $Tr(T)=k+1$ for a $tcmbt_{k}$. ∎

### v2

@@ 2.2 Basic properties of total transitivity
For any graph $G$, $Tr_{t}(G)\leq\min\{\Delta(G),\left\lfloor\frac{n}{2}\right\rfloor\}$, where $\Delta(G)$ is the maximum degree of $G$.
@@
Let $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$ be a $Tr_{t}$-transitive partition of $G$. Let $x\in V_{k}$. As $V_{1}$ is a total dominating set of $G$, and for $11\leq i\leq k-$ 1, $V_{i+1}$ is a total dominating set in the graph $G_{i}=G-\bigcup_{j=1}^{i}V_{j}$, $deg(x)\geq k$ and $V_{j}\geq 2$ for all $1\leq j\leq k$. Therefore, $Tr_{t}(G)=k\leq deg(x)\leq\Delta(G)$ and $Tr_{t}(G)\leq\left\lfloor\frac{n}{2}\right\rfloor$. Hence, $Tr_{t}(G)\leq\min\{\Delta(G),\left\lfloor\frac{n}{2}\right\rfloor\}$. ∎
@@
The partition $\{V_{1},V_{2},\ldots,V_{\left\lfloor\frac{n}{2}\right\rfloor}\}$ such that each set contains at least two vertices is a total transitive partition of $K_{n}$ of size $\left\lfloor\frac{n}{2}\right\rfloor$. Moreover, from the Proposition 1, we have $Tr_{t}(G)\leq\min\{\Delta(K_{n}),\left\lfloor\frac{n}{2}\right\rfloor\}=\min\{n,\left\lfloor\frac{n}{2}\right\rfloor\}=\left\lfloor\frac{n}{2}\right\rfloor$. Therefore, $Tr_{t}(K_{n})=\left\lfloor\frac{n}{2}\right\rfloor$. ∎
@@
$Tr_{t}(P_{n})=\begin{cases}1&n=2,3,4,5\\ 2&n\geq 6\end{cases}$
@@
Let $C_{n}$ be a cycle of $n$ vertices, then the total transitivity of $C_{n}$ is given as follows:
$Tr_{t}(C_{n})=\begin{cases}1&n=3\\ 2&n\geq 4\end{cases}$
@@
For a complete bipartite graph $K_{m,n}$ with $m+n$ vertices $Tr_{t}(K_{m,n})=\min\{m,n\}$.
@@
Let the vertex set of $K_{m,n}$ be $V=X\cup Y$, where $|X|=m$, $|Y|=n$, and assume $m\leq n$. Consider the vertex partition $\pi=\{V_{1},V_{2},\dots,V_{m}\}$, where $V_{i}=\{x_{i},y_{i}\}$ for all $2\leq i\leq m$, and $V_{1}$ contains all remaining vertices of $K_{m,n}$. Since $K_{m,n}$ is a complete bipartite graph, this partition is a total transitive partition of size $m$, implying that $Tr_{t}(K_{m,n})\geq m$.
Next, we show that $Tr_{t}(K_{m,n})$ cannot exceed $m$. Suppose, for contradiction, that $Tr_{t}(K_{m,n})=k\geq m+1$, and let $\pi^{\prime}=\{V_{1},V_{2},\dots,V_{k}\}$ be a total transitive partition of $K_{m,n}$ with size $k$. Choose any vertex $v\in V_{k}$. Since $v$ must be totally dominated within $V_{k}$, it follows that the degree of $v$ must be at least $k\geq m+1$. However, this condition is only possible if $v\in X$, as vertices in $Y$, have a degree at most $m$. Moreover, since $V_{k}$ totally dominates itself, there must exist $u\in V_{k}$ such that $uv\in E$ and $u\in Y$. Since $u\in V_{k}$, it must also have a degree at least $k\geq m+1$. But this contradicts the fact that $\deg(u)=m<k$. Hence, no such partition of size greater than $m$ can exist, implying that $Tr_{t}(K_{m,n})\leq m$. Thus, we conclude that $Tr_{t}(K_{m,n})=\min\{m,n\}$. ∎
@@
If $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$ is a total transitive partition of a graph $G$, then for any two sets $V_{i}$ and $V_{j}$ in $\pi$, where $i<j$, the partition $\pi^{\prime}=\{V_{1},V_{2},\ldots,V_{i-1},V_{i}\cup V_{j},V_{i+1},\ldots,V_{j-1},V_{j+1},\ldots,V_{k}\}$ is a total transitive partition.
Proof.
By the definition of total transitive partition, clearly it follows that $\pi^{\prime}$ is a total transitive partition. of $G$ of size $k-1$. ∎
@@
Let $G$ be a disconnected graph with connected components $C_{1},C_{2},\ldots,C_{t}$. Then $Tr_{t}(G)=\max\{Tr_{t}(C_{i})|1\leq i\leq k\}$.
Proof.
Let $C_{j}$ be the component of $G$ such that $\max\{Tr_{t}(C_{i})|1\leq i\leq k\}=Tr_{t}(C_{j})=p$. Furthermore, assume that $\pi=\{V_{1},V_{2},\ldots,V_{p}\}$ is a total transitive partition of $C_{j}$ of size $p$. Consider $\pi^{\prime}=\{(\bigcup_{r=1,r\neq j}^{t}C_{r})\cup V_{1},V_{2},\ldots,V_{p}\}$. Clearly, $\pi^{\prime}$ is a total transitive partition of $G$ of size $p$ and hence $Tr_{t}(G)\geq p=Tr_{t}(C_{j})=\max\{Tr_{t}(C_{i})|1\leq i\leq k\}$.
For the other part, let $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$ be a $Tr_{t}$-partition of $G$. Furthermore, let for some $l\in\{1,2,\ldots,k\}$, $V(C_{l})\cap V_{k}\not=\phi$. Since $V_{j}$ totally dominates $V_{k}$ for all $1\leq j\leq k$, then $V(C_{l})\cap V_{j}\not=\phi$. For a similar reason, we have $(V_{j}\cap V(C_{l}))$, which totally dominates $(V_{k}\cap V(C_{l}))$. Therefore, $\{V(C_{l})\cap V_{1},V(C_{l})\cap V_{2},\ldots,V(C_{l})\cap V_{k}\}$ is a total transitive partition of $C_{l}$. Thus, $k=Tr_{t}(G)\leq Tr_{t}(C_{l})$, that is, $Tr_{t}(G)\leq\max\{Tr_{t}(C_{i})|1\leq i\leq k\}$. Hence, $Tr_{t}(G)=\max\{Tr_{t}(C_{i})|1\leq i\leq k\}$. ∎
Now we show that if the total transitivity of a graph $G$ is constant, that is, independent of the number of vertices, and $\pi=\{V_{1},V_{2},\ldots,V_{k-1},V_{k}\}$ is a total transitive partition of $G$, then the number of vertices of each $V_{i}$, for $2\leq i\leq k$, does not depend on $n$; rather, it is bounded by some exponential in $i$.
Proposition 9.
@@
Let $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$ be a total transitive partition of $G$ of size $k$, where $k\geq 3$. Let $x\in V_{k}$; as $V_{k}$ totally dominates $V_{k}$, there exists $y\in V_{k}$ such that $xy\in E(G)$. So, $|V_{k}|\geq 2$. Now if $|V_{k}-\{x,y\}|\geq 1$, then we move all the vertices except two, say vertices $x$ and $y$, from the set $V_{k}$ to the first set $V_{1}$. After this transformation, we have $\pi^{\prime}=\{V_{1}^{\prime}=V_{1}\cup(V_{k}-\{x,y\}),V_{2},\ldots,V_{k-1},V_{k}^{\prime}=\{x,y\}\}$, which is a total transitive partition of $G$ of size $k$ and $|V_{k}|=2$. Therefore, $\pi^{\prime}$ is a total transitive of $G$ of size $k$, such that $|V_{k}|=2$. Consider the base case when $i=1$. Note that to totally dominate $V_{k}$ and $V_{k-1}$, we need at most $2\cdot|V_{k}|$ vertices in $V_{k-1}$, and the remaining vertices can be moved to $V_{1}$ as in the previous argument. Therefore, $|V_{k-1}|\leq 2\cdot|V_{k}|=2\cdot 2=4=2^{i+1}$. Assume the induction hypothesis and consider that we have a total transitive partition $\pi=\{V_{1},V_{2},\ldots\,V_{k}\}$, such that $|V_{k}|=2$ and $|V_{k-j}|\leq 2^{j+1}$, $11\leq j\leq i-$ 1. Now consider the set $V_{k-i}$. Similarly, at most $2\cdot|V_{k-i+1}|$ vertices in $V_{k-i}$ are needed to totally dominate the $i+1$ sets that follow and include it. If $V_{k-i}$ contains more than that many vertices, the remaining vertices can be moved to $V_{1}$, maintaining the total transitivity of $G$. Hence, $V_{k-i}\leq 2\cdot|V_{k-i+1}|=2\cdot 2^{i}=2^{i+1}$. ∎
In the Remark 1, we have already shown that the difference between transitivity and total transitivity can be arbitrarily large. Now, we show another non-trivial example where the difference between these two parameters is arbitrarily large. Let $G=(S\cup K,E)$ be a split graph, where $S$ and $K$ are an independent set and clique of $G$, respectively. Furthermore, assume that $K$ is the maximum clique of $G$, that is, $\omega(G)=|K|$. Furthermore, $|K|=|S|$, and each vertex of $S$ is adjacent to exactly one vertex of $K$ (see Figure 1).
@@
First, we show that for the split graph given in Figure 1, $Tr_{t}(G)=1$. Let $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$ be a total transitive partition of the above split graph $G$. Since $deg(s)=1$ for all $s\in S$, $s$ cannot be in $V_{p}$ for $p\geq 2$. Moreover, $\pi$ is a total transitive partition, which implies that to totally dominate $s\in S$, there must be a vertex from $K$ in $V_{1}$. From the above example graph, it is clear that to totally dominate vertices from $S$, all the vertices of $K$ must be in $V_{1}$. Hence, $Tr_{t}(G)=1$. Again, we know from [19], $Tr(G)=\omega(G)+1$, as $S$ dominates $K$. Therefore, for large $\omega(G)$, the difference $Tr(G)-Tr_{t}(G)=\omega(G)+1-1=\omega(G)$ is arbitrarily large.
Next, we prove the existence of graphs with transitivity is one more than total transitivity. We define a special tree, namely total transitive tree of order $k$ and we denote it $\mathcal{T}_{k}$. Apart from this property, the $\mathcal{T}_{k}$ possesses some important properties. The recursive definition of $\mathcal{T}_{k}$ is as follows:
@@
The total transitive tree of order $k$, denoted as $\mathcal{T}_{k}$, is defined recursively as follows:
@@
    The total transitive tree of order $1$, denoted as $\mathcal{T}_{1}$, is a rooted tree, whose root has degree $1$ and adjacent to a vertex of degree $1$.
@@
    The total transitive tree of order $k$, denoted as $\mathcal{T}_{k}$, is a rooted tree whose root has degree $k$ and is adjacent to vertices $\{v_{1},v_{2},\ldots,v_{k}\}$. For all $11\leq i\leq k-$ 1, $v_{i}$ are roots of the total transitive tree of order $i$, denoted by $\mathcal{T}_{i}$. Moreover, the vertex $v_{k}$ is adjacent to vertices $\{u_{1},u_{2},\ldots,u_{k-1}\}$, such that $u_{i}$ are roots of the total transitive tree of order $i$ ($\mathcal{T}_{i}$), for all $11\leq i\leq k-$ 1.
Figure 2: Total transitive tree of order up to $3$ along with its unique total transitive partition
Figure 2 illustrates the total transitive tree of order up to $3$. Note that from the definition, it follows that the root of $\mathcal{T}_{k}$ and exactly one of its children are the only vertices that can be in $V_{k}$, in a total transitive partition of order $k$. Because of the acyclic property, none of the vertices in $V_{i}$ can dominate more than one vertex from $V_{j}$ with $j>i$. So, $\mathcal{T}_{k}$ is the smallest tree (in terms of vertices) with total transitivity equal to $k$. It also has the only total transitive partition of size $k$ (see Figure 2). This property of $\mathcal{T}_{k}$ helps us to use these graphs as a gadget in the NP-completeness proofs in Section 4. We show that if a graph $G$ is a $\mathcal{T}_{k}$, then $Tr_{t}(G)=k$ and $Tr(G)=k+1$. For that, we have the following lemma.
@@
First, we show that $Tr_{t}(T)=k$. For this, we use induction on the order of the total transitive tree. When $k=1$, the base case, $T$, is a $P_{2}$ by the definition of a total transitive tree. Therefore, $Tr_{t}(P_{2})=1$. Let $T$ be a $\mathcal{T}_{k}$. By the definition, we have $T$ is a rooted tree, whose root has degree $k$ and is adjacent to vertices $\{v_{1},v_{2},\ldots,v_{k}\}$, such that $v_{i}$ are roots of the total transitive tree of order $i$ $(\mathcal{T}_{i})$, for all $11\leq i\leq k-$ 1. Moreover, the vertex $v_{k}$ is adjacent to vertices $\{u_{1},u_{2},\ldots,u_{k-1}\}$, such that $u_{i}$ are roots of the total transitive tree of order $i$ $(\mathcal{T}_{i})$, for all $11\leq i\leq k-$ 1. Let $c$ be the root of $T$ and $c^{\prime}$ be its child such that $c^{\prime}$ is adjacent to vertices $\{u_{1},u_{2},\ldots,u_{k-1}\}$, such that $u_{i}$ are roots of the total transitive tree of order $i$ $(\mathcal{T}_{i})$, for all $11\leq i\leq k-$ 1. By the induction hypothesis, $Tr_{t}(T_{v_{j}})=Tr_{t}(T_{u_{j}})=j$, for all $11\leq i\leq k-$ 1, where $T_{v}$ is a subtree of $T-\{c\}$, rooted at $v$ (see the $\mathcal{T}_{3}$ in Figure 2). Now $T-\{c,c^{\prime}\}$ is a disjoint union of trees $T_{v_{j}},T_{u_{j}}$, for all $11\leq j\leq k-$ 1, and $k-1=\max\{Tr_{t}(T_{v_{j}})|1\leq j\leq k-1\}=\max\{Tr_{t}(T_{u_{j}})|1\leq j\leq k-1\}$. Total transitivity depends only on the maximum total transitivity of a component of a disconnected graph (Proposition 8). Therefore, $Tr_{t}(T-\{c,c^{\prime}\})$ is $k-1$. Moreover, we can construct a total transitive partition, say $\pi=\{V_{1},V_{2},\ldots,V_{k-1}\}$, of $T-\{c,c^{\prime}\}$ recursively such that $\{v_{i},u_{i}\}\in V_{i}$ for all $11\leq i\leq k-$ 1. Now clearly, $\pi^{\prime}=\pi\cup\{c,c^{\prime}\}$ forms a total transitive partition of $T$. Therefore, $Tr_{t}(T)\geq k$. Moreover, by the Proposition 1, we know that $Tr_{t}(T)\leq\Delta(G)=k$. Hence, $Tr_{t}(T)=k$. The algorithm described in [14] shows that $Tr(T)=k+1$. Therefore, $Tr_{t}(T)=k$ and $Tr(T)=k+1$ for $\mathcal{T}_{k}$. ∎